Calculate the mass of X required to prepare 250 mL standard solution of molarity 0.017 moles/dm3
1
Expert's answer
2020-06-03T13:34:00-0400

C(x) = m(x)/(M(x)*V(x)

С(x) = n(x)*V(x)

n(x) = C(x)/V(x)=0.017 moles/dm3 / 0.25 dm3 = 0.068 moles


What is molar mass of X?